# Enhancing Uncertainty Estimation in LLMs with Expectation of Aggregated Internal Belief

### v1

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success across a wide range of natural language tasks, but often exhibit overconfidence and generate plausible yet incorrect answers. This overconfidence, especially in models undergone Reinforcement Learning from Human Feedback (RLHF), poses significant challenges for reliable uncertainty estimation and safe deployment. In this paper, we propose EAGLE (Expectation of AGgregated internaL bEief), a novel self-evaluation-based calibration method that leverages the internal hidden states of LLMs to derive more accurate confidence scores. Instead of relying on the model’s final output, our approach extracts internal beliefs from multiple intermediate layers during self-evaluation. By aggregating these layer-wise beliefs and calculating the expectation over the resulting confidence score distribution, EAGLE produces a refined confidence score that more faithfully reflects the model’s internal certainty. Extensive experiments on diverse datasets and LLMs demonstrate that EAGLE significantly improves calibration performance over existing baselines. We also provide an in-depth analysis of EAGLE, including a layer-wise examination of uncertainty patterns, a study of the impact of self-evaluation prompts, and an analysis of the effect of self-evaluation score range.  [^1] Our code is public at https://github.com/sustech-nlp/EAGLE.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide spectrum of natural language understanding and generation tasks (Brown et al. 2020; Wolf et al. 2020; Achiam et al. 2023; Touvron et al. 2023). Their ability to generate coherent and contextually relevant text has led to their increasing deployment in various real-world applications. However, a critical challenge remains: LLMs often produce plausible-sounding but incorrect answers, a phenomenon known as hallucination (Bai et al. 2022). This challenge is further exacerbated by the lack of reliable uncertainty estimation for the outputs of LLMs that have undergone Reinforcement Learning from Human Feedback (RLHF) (Kadavath et al. 2022; Tian et al. 2023), undermining their trustworthiness and limits their safe deployment. To mitigate this issue, recent research has introduced calibration techniques, which aim to ensure that LLMs’ confidence of predictions accurately reflect their likelihood of correctness (Kadavath et al. 2022; Chen and Mueller 2023; Xiong et al. 2023).
Among the various research directions, a particularly important approach is leveraging the LLMs’ inherent capabilities for self-reflection and self-evaluation as a means of uncertainty estimation (Lin, Hilton, and Evans 2022; Kadavath et al. 2022; Tian et al. 2023; Chen and Mueller 2023; Huang et al. 2024). These methods often involve prompting the LLM to explicitly assess the uncertainty and/or correctness of its own generated responses, usually by eliciting verbalized confidence scores.
The motivation of our method stems from the observation that LLMs’ internal representations, particularly the hidden states at different layers, contain rich information about LLMs’ internal belief and can be used to reveal truthfulness of statements, knowledge discovery and behavior monitoring (Burns et al. 2022; Azaria and Mitchell 2023; Zou et al. 2023). Inspired by this, we propose a novel self-evaluation-based calibration method, EAGLE (Hidden Layer Aggregation), for uncertainty estimation. Unlike previous methods that rely on superficial verbalized confidence scores, our approach derives the confidence score by leveraging the internal representations of LLMs.
@@ 2 Related Work
Model calibration—the alignment of a model’s predicted probabilities with the true likelihood of correctness (Hendrycks and Gimpel 2016; Guo et al. 2017)—is crucial for user trust and informed decision-making in AI systems. In traditional supervised learning, post-hoc calibration techniques such as temperature scaling (Guo et al. 2017), Platt scaling (Platt et al. 1999), and isotonic regression (Niculescu-Mizil and Caruana 2005) are used to improve the reliability of model confidence estimates. However, directly applying these methods to LLMs, particularly those undergone RLHF (Kadavath et al. 2022; Tian et al. 2023), poses significant challenges.
Calibration in LLMs has been approached primarily through two lines of research. The first involves analyzing the diversity of multiple generated outputs, where techniques such as clustering-based metrics are used to examine the variance in the model’s responses (Kuhn, Gal, and Farquhar 2023; Lin, Trivedi, and Sun 2024). While effective in some scenarios, these methods are often computationally intensive due to the need for multiple generations per input, and the diversity observed does not always reliably indicate the correctness of individual outputs. A second, increasingly prominent direction leverages the self-evaluation capabilities of LLMs. In this approach, the model is prompted to assess the quality or correctness of its own generated responses, typically by providing a verbalized confidence score or categorical judgment (Lin, Hilton, and Evans 2022; Kadavath et al. 2022; Tian et al. 2023; Chen and Mueller 2023; Huang et al. 2024). Self-evaluation methods offer a more direct means of accessing the model’s internal assessment of its outputs, and have shown promise for improving calibration. However, most existing self-evaluation approaches rely on superficial verbalized confidence scores, which may fail to capture the nuanced confidence signals present in the model’s internal representations.
@@
Analyzing the internal representations of LLMs has emerged as a crucial direction for understanding their inner workings and decision-making processes. A promising line of research involves interpreting these hidden states by projecting them back into the vocabulary space. This technique allows researchers to treat the representation at each layer as a distribution over the vocabulary, offering insights into the model’s predictions and internal ”beliefs” at different stages of processing (Nostalgebraist 2020; Geva et al. 2022). Recent studies have leveraged this approach to probe the model’s internal states for various purposes, such as discovering latent knowledge (Burns et al. 2022), assessing the truthfulness of statements (Azaria and Mitchell 2023), and even steering model behavior through representation engineering (Zou et al. 2023). These works collectively demonstrate that a model’s internal representations often contain more nuanced and faithful information than its final output layer alone.
@@ 4 Methodology
In this work, we focus on the problem of uncertainty estimation for LLMs, which is particularly challenging for LLMs that have undergone RLHF (Kadavath et al. 2022; Tian et al. 2023). A key concept in uncertainty estimation is calibration. A model is considered well-calibrated if its confidence of predictions accurately reflect their likelihood of correctness. Formally, for a given input $x$, its golden label $y$, and model prediction $\hat{y}$, a perfectly calibrated model satisfies:
@@
The most commonly used metric for quantifing the degree of miscalibration is Expected Calibration Error (ECE). ECE is defined as $\mathbb{E}[|\operatorname{Pr}(\hat{y}=y\mid C(\hat{y}|x)=p)-p|]$, which measures the difference between confidence and accuracy. An empirical estimate of ECE is calculated by partitioning $N$ samples into $M$ bins $\{B_{1},B_{2},\ldots,B_{M}\}$ according to the confidence score $C$ predicted by the model. The ECE is then formulated as:
@@
Our motivation stems from a growing body of work demonstrating that the internal representations of LLMs, specifically the hidden states across different layers, encode a wealth of information that is not always reflected in the final output. Research has shown that these internal states can be probed to discover latent knowledge, track the truthfulness of statements, and monitor the model’s behavior (Burns et al. 2022; Azaria and Mitchell 2023; Zou et al. 2023). For instance, Burns et al. (2022) found that a simple probe applied to hidden states could consistently reveal a model’s latent knowledge, even when the model’s final output was incorrect. Similarly, Zou et al. (2023) demonstrated that representations of concepts could be identified and steered within the model. This suggests that a model’s true “belief” or uncertainty about its own answer is more accurately represented in the collective information distributed across its layers, rather than in a single, potentially biased, verbalization. Therefore, we hypothesize that by aggregating information directly from the hidden states associated with a self-evaluation token, we can construct a more robust and better-calibrated uncertainty estimate.
Furthermore, we posit that relying on a single point estimate for the confidence score, such as selecting the most probable score, may fail to capture the full extent of the model’s uncertainty. The entire probability distribution over possible scores contains richer information about the model’s internal belief state. Based on these insights, our approach, which we call EAGLE, taps into the deeper, multi-layered internal reasoning of the model, bypassing the final layer’s potential for overconfident posturing. It then calculates the expected value over the entire confidence score distribution to capture a more nuanced and faithful measure of the model’s internal belief.
EAGLE: Expectation of AGgregated internaL bEief
Given a question $x$ and the LLM’s generated answer $\hat{y}$, we prompt the LLM to evaluate its own answer. The self-evaluation prompt is designed to elicit a confidence score about the correctness of the answer:
@@
Rather than merely selecting the confidence score with the highest probability like previous works (Lin, Hilton, and Evans 2022; Kadavath et al. 2022; Tian et al. 2023; Chen and Mueller 2023; Huang et al. 2024), we utilize the entire distribution to derive a more robust estimate of the model’s confidence. This approach allows us to capture the nuanced uncertainty in the model’s self-evaluation, avoiding reliance on a single point estimate that may not fully represent the model’s internal belief. Formally, the final confidence score, $C(\hat{y}|x)$, is calculated as the examination of the confidence score distribution:
@@ 5 Experiments
We evaluate the calibration performance of various methods on both open-ended and multiple-choice datasets. For open-ended tasks, we use rc.nocontext subset of TriviaQA (Joshi et al. 2017) and GSM8k (Cobbe et al. 2021). For multiple-choice tasks, we include MMLU (Hendrycks et al. 2021). The final evaluation is conducted on the TriviaQA validation set, GSM8k test set, and MMLU test set. For hyperparameter tuning, we use an equal-size subset of the TriviaQA and GSM8k training sets, and the official validation set for MMLU. These datasets cover diverse domains and reasoning complexities, enabling a comprehensive assessment of calibration across different tasks.
@@
We evaluate our method on two prominent families of LLMs: Qwen2.5 (Team 2024) and Llama3 (AI@Meta 2024). Specifically, we include Qwen2.5 models of varying scales (7B and 72B) and Llama3 models (8B and 70B). These models are selected to cover a wide range of parameter sizes and architectures, enabling a comprehensive assessment of calibration performance across different model capacities. For clarity, we refer to Qwen2.5-7B-Instruct and Meta-Llama-3-8B-Instruct as Qwen2.5 7B and Llama3 8B, respectively, and adopt similar abbreviations for other model variants throughout the paper.
@@
  • Semantic Entropy (SE) (Kuhn, Gal, and Farquhar 2023): An entropy measure that clusters semantically equivalent answers before computing uncertainty, addressing linguistic variations of correct responses.
  • Self-consistency (Self-cons.) (Chen and Mueller 2024): Confidence derived from agreement among multiple sampled answers, with higher agreement indicating greater certainty.
  • P(True) (Kadavath et al. 2022): The model’s self-evaluation of whether its answer is correct, obtained by prompting it to classify answers as true/false.
  • Contextualized Sequence Likelihood (CSL) (Lin, Trivedi, and Sun 2024): A weighted sequence likelihood that uses attention mechanisms to focus on relevant tokens when computing answer probabilities.
@@
Table 1 presents the main experimental results, comparing the calibration performance of our proposed method, EAGLE, against several baseline. The results demonstrate that EAGLE consistently achieves superior calibration performance.
@@
In summary, the empirical results strongly support our hypothesis. By aggregating internal beliefs from multiple layers, EAGLE provides a more faithful and reliable measure of an LLM’s uncertainty. Its consistent and significant improvements over strong baselines across different models, scales, and tasks underscore its potential as a general-purpose calibration method.
@@
In summary, the ablation study confirms that each component of the EAGLE framework is crucial for its effectiveness. The optimal configuration—aggregating logits from the last-$n$ layers and calculating the expectation of the resulting score distribution—consistently delivers the best calibration performance across all tested models and datasets.
6 Analysis
@@
Prompt Variant A (Without score criteria)
Question: [QUESTION] Answer: [LLM’S ANSWER] Rate how likely the answer is to be correct using a number from 0 to 9: Confidence score (0-9):
Prompt Variant B (CoT)
Question: [QUESTION] Answer: [LLM’S ANSWER] Rate how likely the answer is to be correct using a number from 0 to 9: - 0 = Very likely incorrect - 5 = Uncertain or partially correct - 9 = Very likely correct You should first provide an explanation in a few sentences, then provide your score. The output format should look as follows: Explanation: <write your explanation here>Confidence score (0-9): <give your score here>
Table 4: Two variants of self-evaluation prompts used to study the sensitivity of EAGLE to prompt formulation.
@@
To investigate the robustness of EAGLE to different self-evaluation prompt formulations, we evaluate the calibration performance using two variants of self-evaluation prompts. The details of these prompt variants are shown in Table 4.
The results in Table 5 show that our method exhibits a robustness to the prompt, although the choice of prompt does have an impact on performance. Our main prompt, which provides clear criteria for the scores, generally yields the best or highly competitive results. For instance, on the GSM8K dataset, all three prompt variants produce similar ECE and AUROC scores for both models, indicating that EAGLE is not very sensitive in this scenario. However, on TriviaQA, the performance varies more. The CoT prompt (Variant B) leads to a noticeable degradation in calibration for both models. This suggests that while CoT can encourage reasoning, it may also introduce noise into the self-evaluation process, which can interfere with the internal confidence signals that EAGLE relies on. Overall, while the performance is not entirely independent of the prompt design, EAGLE demonstrates reasonable stability, particularly when provided with a clear and direct self-evaluation format.
Effect of Confidence Score Range

### v2

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success across a wide range of natural language tasks, but often exhibit overconfidence and generate plausible yet incorrect answers. This overconfidence, especially in models undergone Reinforcement Learning from Human Feedback (RLHF), poses significant challenges for reliable uncertainty estimation and safe deployment. In this paper, we propose EAGLE (Expectation of AGgregated internaL bEief), a novel self-evaluation-based calibration method that leverages the internal hidden states of LLMs to derive more accurate confidence scores. Instead of relying on the model’s final output, our approach extracts internal beliefs from multiple intermediate layers during self-evaluation. By aggregating these layer-wise beliefs and calculating the expectation over the resulting confidence score distribution, EAGLE produces a refined confidence score that more faithfully reflects the model’s internal certainty. Extensive experiments on diverse datasets and LLMs demonstrate that EAGLE significantly improves calibration performance over existing baselines. We also provide an in-depth analysis of EAGLE, including a layer-wise examination of uncertainty patterns, a study of the impact of self-evaluation prompts, and an analysis of the effect of self-evaluation score range. Our code is public at https://github.com/sustech-nlp/EAGLE.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide spectrum of natural language understanding and generation tasks (Brown2020LanguageMA; wolf2020huggingfacestransformersstateoftheartnatural; Achiam2023GPT4TR; Touvron2023LLaMAOA). Their ability to generate coherent and contextually relevant text has led to their increasing deployment in various real-world applications. However, a critical challenge remains: LLMs often produce plausible-sounding but incorrect answers, a phenomenon known as hallucination (Bai2022ConstitutionalAH). This challenge is further exacerbated by the lack of reliable uncertainty estimation for the outputs of LLMs that have undergone Reinforcement Learning from Human Feedback (RLHF) (kadavath2022language; Tian2023JustAF), undermining their trustworthiness and limits their safe deployment. To mitigate this issue, recent research has introduced calibration techniques, which aim to ensure that LLMs’ confidence of predictions accurately reflect their likelihood of correctness (kadavath2022language; Chen2023QuantifyingUI; xiong2023can).
Among the various research directions, a particularly important approach is leveraging the LLMs’ inherent capabilities for self-reflection and self-evaluation as a means of uncertainty estimation (Lin2022TeachingMT; kadavath2022language; Tian2023JustAF; Chen2023QuantifyingUI; Huang2024CalibratingLG). These methods often involve prompting the LLM to explicitly assess the uncertainty and/or correctness of its own generated responses, usually by eliciting verbalized confidence scores.
The motivation of our method stems from the observation that LLMs’ internal representations, particularly the hidden states at different layers, contain rich information about LLMs’ internal belief and can be used to reveal truthfulness of statements, knowledge discovery and behavior monitoring (burns2022discovering; azaria2023internal; zou2023representation; laibeyond). In particular, concurrent work (ji2025calibrating) demonstrates that intermediate-layer hidden states naturally separate high-confidence from low-confidence model predictions, suggesting that the model’s latent space already encodes information about its own uncertainty. Inspired by this, we propose a novel self-evaluation-based calibration method, EAGLE (Hidden Layer Aggregation), for uncertainty estimation. Unlike previous methods that rely on superficial verbalized confidence scores, our approach derives the confidence score by leveraging the internal representations of LLMs.
@@ 2 Related Work
Model calibration—the alignment of a model’s predicted probabilities with the true likelihood of correctness (Hendrycks2016ABF; Guo2017OnCO)—is crucial for user trust and informed decision-making in AI systems. In traditional supervised learning, post-hoc calibration techniques such as temperature scaling (Guo2017OnCO), Platt scaling (platt1999probabilistic), and isotonic regression (niculescu2005predicting) are used to improve the reliability of model confidence estimates. However, directly applying these methods to LLMs, particularly those undergone RLHF (kadavath2022language; Tian2023JustAF), poses significant challenges.
Calibration in LLMs has been approached primarily through two lines of research. The first involves analyzing the diversity of multiple generated outputs, where techniques such as clustering-based metrics are used to examine the variance in the model’s responses (kuhn2023semantic; lin2024contextualized). While effective in some scenarios, these methods are often computationally intensive due to the need for multiple generations per input, and the diversity observed does not always reliably indicate the correctness of individual outputs. A second, increasingly prominent direction leverages the self-evaluation capabilities of LLMs. In this approach, the model is prompted to assess the quality or correctness of its own generated responses, typically by providing a verbalized confidence score or categorical judgment (Lin2022TeachingMT; kadavath2022language; Tian2023JustAF; Chen2023QuantifyingUI; Huang2024CalibratingLG). Self-evaluation methods offer a more direct means of accessing the model’s internal assessment of its outputs, and have shown promise for improving calibration. However, most existing self-evaluation approaches rely on superficial verbalized confidence scores, which may fail to capture the nuanced confidence signals present in the model’s internal representations.
@@
Analyzing the internal representations of LLMs has emerged as a crucial direction for understanding their inner workings and decision-making processes. A promising line of research involves interpreting these hidden states by projecting them back into the vocabulary space. This technique allows researchers to treat the representation at each layer as a distribution over the vocabulary, offering insights into the model’s predictions and internal ”beliefs” at different stages of processing (nostalgebraist2020interpreting; Geva2022TransformerFL). Recent studies have leveraged this approach to probe the model’s internal states for various purposes, such as discovering latent knowledge (burns2022discovering), assessing the truthfulness of statements (azaria2023internal), and even steering model behavior through representation engineering (zou2023representation). These works collectively demonstrate that a model’s internal representations often contain more nuanced and faithful information than its final output layer alone.
@@ 4 Methodology
In this work, we focus on the problem of uncertainty estimation for LLMs, which is particularly challenging for LLMs that have undergone RLHF (kadavath2022language; Tian2023JustAF). A key concept in uncertainty estimation is calibration. A model is considered well-calibrated if its confidence of predictions accurately reflect their likelihood of correctness. Formally, for a given input $x$, its golden label $y$, and model prediction $\hat{y}$, a perfectly calibrated model satisfies:
@@
The most commonly used metric for quantifying the degree of miscalibration is Expected Calibration Error (ECE). ECE is defined as $\mathbb{E}[|\operatorname{Pr}(\hat{y}=y\mid C(\hat{y}|x)=p)-p|]$, which measures the difference between confidence and accuracy. An empirical estimate of ECE is calculated by partitioning $N$ samples into $M$ bins $\{B_{1},B_{2},\ldots,B_{M}\}$ according to the confidence score $C$ predicted by the model. The ECE is then formulated as:
@@
Our motivation stems from a growing body of work demonstrating that the internal representations of LLMs, specifically the hidden states across different layers, encode a wealth of information that is not always reflected in the final output. Research has shown that these internal states can be probed to discover latent knowledge, track the truthfulness of statements, and monitor the model’s behavior (burns2022discovering; azaria2023internal; zou2023representation; laibeyond). For instance, burns2022discovering found that a simple probe applied to hidden states could consistently reveal a model’s latent knowledge, even when the model’s final output was incorrect. Similarly, zou2023representation demonstrated that representations of concepts could be identified and steered within the model. In particular, concurrent work (ji2025calibrating) demonstrates that intermediate layer hidden states can separate high-confidence from low-confidence model predictions, suggesting that the model’s latent space already encodes information about its own uncertainty.
Inspired by these empirical findings, we hypothesize that by aggregating information directly from hidden states of intermediate layers, we can construct a more robust and better-calibrated uncertainty estimate. Furthermore, we posit that relying on a single point estimate for the confidence score, such as selecting the most probable score, may fail to capture the full extent of the model’s uncertainty, as the entire probability distribution over possible scores contains richer information about the model’s internal belief state. Based on these insights, our approach, which we call EAGLE, taps into the deeper, multi-layered internal reasoning of the model, bypassing the final layer’s potential for overconfident posturing. It then uses the expectation of the confidence score distribution to capture a more nuanced and faithful measure of the model’s internal belief.
EAGLE: Expectation of Aggregated Internal Belief
Given a question $x$ and the LLM’s generated answer $\hat{y}$, we prompt the LLM to evaluate its own answer. The self-evaluation prompt is designed to elicit a confidence score about the correctness of the answer.
[…]
@@
Rather than merely selecting the confidence score with the highest probability like previous works (Lin2022TeachingMT; kadavath2022language; Tian2023JustAF; Chen2023QuantifyingUI; Huang2024CalibratingLG), we utilize the entire distribution to derive a more robust estimate of the model’s confidence. This approach allows us to capture the nuanced uncertainty in the model’s self-evaluation, avoiding reliance on a single point estimate that may not fully represent the model’s internal belief. Formally, the final confidence score, $C(\hat{y}|x)$, is calculated as the examination of the confidence score distribution:
@@ 5 Experiments
We evaluate the calibration performance of various methods on both open-ended and multiple-choice datasets. For open-ended tasks, we use rc.nocontext subset of TriviaQA (joshi2017triviaqa) and GSM8k (cobbe2021trainingverifierssolvemath). For multiple-choice tasks, we include MMLU (hendryckstest2021). The final evaluation is conducted on the TriviaQA validation set, GSM8k test set, and MMLU test set. For hyperparameter tuning, we use an equal-size subset of the TriviaQA and GSM8k training sets, and the official validation set for MMLU. These datasets cover diverse domains and reasoning complexities, enabling a comprehensive assessment of calibration across different tasks.
@@
We evaluate our method on two prominent families of LLMs: Qwen2.5 (qwen2.5) and Llama3 (grattafiori2024llama). Specifically, we include Qwen2.5 models of varying scales (7B and 72B) and Llama3 models (8B and 70B). These models are selected to cover a wide range of parameter sizes and architectures, enabling a comprehensive assessment of calibration performance across different model capacities. For clarity, we refer to Qwen2.5-7B-Instruct and Meta-Llama-3-8B-Instruct as Qwen2.5 7B and Llama3 8B, respectively, and adopt similar abbreviations for other model variants throughout the paper.
@@
  • Semantic Entropy (SE) (kuhn2023semantic): An entropy measure that clusters semantically equivalent answers before computing uncertainty, addressing linguistic variations of correct responses.
  • Self-consistency (Self-cons.) (chen2024quantifying): Confidence derived from agreement among multiple sampled answers, with higher agreement indicating greater certainty.
  • P(True) (kadavath2022language): The model’s self-evaluation of whether its answer is correct, obtained by prompting it to classify answers as true/false.
  • Contextualized Sequence Likelihood (CSL) (lin2024contextualized): A weighted sequence likelihood that uses attention mechanisms to focus on relevant tokens when computing answer probabilities.
@@
Table 1 presents the main experimental results, comparing the calibration performance of our proposed method, EAGLE, against several baselines [^1] We observe that our reported CSL results differ from those presented in the original paper (lin2024contextualized). The discrepancy arises from the use of different model types: the original CSL results were obtained using base LLMs, whereas our experiments employ instruction-tuned (RLHF) models. RLHF models are known to exhibit weaker calibration compared to base LLMs (Tian2023JustAF), leading to the observed divergence.. The results demonstrate that EAGLE consistently achieves superior calibration performance.
@@
6 Analyses
@@
To investigate the robustness of EAGLE to different self-evaluation prompt formulations, we evaluate the calibration performance using two variants of self-evaluation prompts. The details of these prompt variants are shown below:
Prompt Variant A (Without score criteria) Question: [QUESTION] Answer: [LLM’S ANSWER] Rate how likely the answer is to be correct using a number from 0 to 9: Confidence score (0-9):
Prompt Variant B (CoT) Question: [QUESTION] Answer: [LLM’S ANSWER] Rate how likely the answer is to be correct using a number from 0 to 9: - 0 = Very likely incorrect - 5 = Uncertain or partially correct - 9 = Very likely correct You should first provide an explanation in a few sentences, then provide your score. The output format should look as follows: Explanation: <write your explanation here> Confidence score (0-9): <give your score here>
The results in Table 4 show that our method exhibits robustness to the prompt, although the choice of prompt does have an impact on performance. Our main prompt, which provides clear criteria for the scores, generally yields the best or highly competitive results. For instance, on the GSM8K dataset, all three prompt variants produce similar ECE and AUROC scores for both models, indicating that EAGLE is not very sensitive in this scenario. However, on TriviaQA, the performance varies more. The CoT prompt (Variant B) leads to a noticeable degradation in calibration for both models. This suggests that while CoT can encourage reasoning, it may also introduce noise into the self-evaluation process, which can interfere with the internal confidence signals that EAGLE relies on. Overall, while the performance is not entirely independent of the prompt design, EAGLE demonstrates reasonable stability, particularly when provided with a clear and direct self-evaluation format.
Effect of Confidence Score Range